OpenStudy (anonymous):

If someone would help me, that would be great. Find the exact value of the radical expression in simplest form. the square root of 169 − the square root of 50 − the square root of 8 13 − 7 the square root of 2 13 − 3 the square root of 2 7 the square root of 2 6 Find the exact value of the radical expression in simplest form. r times the square root of r − 3 times r times the square root of r + 4 times r times the square root of r 2r 2 times r times the square root of r 2 the square root of r the square root of r

OpenStudy (anonymous):

first off \(\sqrt{169}=13\) right?

OpenStudy (anonymous):

and since \(50=25\times 2\)we have \(\sqrt{50}=\sqrt{25\times 2}=\sqrt{25}\sqrt{2}=5\sqrt{2}\)

OpenStudy (anonymous):

similarly \(\sqrt{8}=\sqrt{4}\sqrt{2}=2\sqrt{2}\)

OpenStudy (anonymous):

so the first one, you can start with \[13-5\sqrt2-2\sqrt2\] and then combine like terms